OpenStudy (anonymous):

Intergral of (x^4/(1+x^6))^2 dx over -1,inf

OpenStudy (anonymous):

Do I read this right? \[ \large \int_{-1}^\infty \frac{x^8}{(1+x^6)^2} dx\]

OpenStudy (anonymous):

Partial fraction decomposition and a lot of spare time (-:

OpenStudy (anonymous):

Can u help me I have 12 minutes left on a test

OpenStudy (anonymous):

I doubt that this problem can be solved within 12 minutes, except there is some fundamental trick I am missing here or I am unable to see, anyone any ideas?

OpenStudy (anonymous):

if the denominator was something in the form x^5, it would be possible, but it's not.

OpenStudy (anonymous):

except you copied the problem wrong, this one is pretty hairy.

OpenStudy (anonymous):

if the numerator*

OpenStudy (anonymous):

Ok what about this one.....

OpenStudy (anonymous):

Integral of 1/(4 sqrt(1+x)) dx over 0,inf

OpenStudy (anonymous):

\[ \large \frac{1}{4} \int_0^\infty (1+x)^{-\frac{1}{2}}dx \]

OpenStudy (anonymous):

Continue

OpenStudy (anonymous):

6minutes left

OpenStudy (anonymous):

oops sorry. well then integrate it.

OpenStudy (anonymous):

\[ \large \frac{1}{2} \sqrt{1+x} \]

OpenStudy (anonymous):

and this integral does not converge obviously, hence there is no limit.

OpenStudy (anonymous):

and that's your answer.

OpenStudy (anonymous):

U can continue I have extra time

OpenStudy (anonymous):

if you want.

OpenStudy (anonymous):

Yes plz

OpenStudy (anonymous):

But this problem is already finished, there isn't anything more to add.

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

\[ \large \left. \frac{1}{2}\sqrt{1+x} \right|_0^\infty \] = does not converge, there is no solution.